A research study assessing vital signs for 15 clients has the following results. Calculate the mean, median, mode for each of the vital sign indices. Client Number Oral Temperature (°F) Pulse (Beats/Minute) Respirations (Breaths/Minute) Blood Pressure  1 97.6 80 12 160/80  2 98.6 60 20 154/90  3 98.6 54 32 132/60  4 99.0 92 16 90/62  5 98.0 86 18 200/140  6 98.4 84 22 116/76  7 99.2 74 28 132/80  8 100.0 72 18 124/78  9 98.6 90 20 140/90 10 98.6 88 32 160/90 11 98.6 64 30 100/50 12 98.8 68 20 118/84 13 98.4 74 18 120/88 14 98.2 50 14 132/74 15 97.6 100 32 190/110
Answer #1

For calculating Mean its basic formula is (sum of the value / total number of the value)

The median is the middle number of the given series after arranged in ascending order

The mode is the most repetitive value of the given series of numbers.

For Temperature;

mean ; 1478.2 / 15 = 98.54
median = 98.6
mode = 98.6


For pulse:

mean; 1136 / 15 = 75.73
median = 74
mode = 74


For respiration
mean; 332 / 15 = 22.13
median = 20
mode = 18,20,32


For BP
mean; 24.6 / 15 = 1.64
median = 1.6
mode = 1.7